OpenStudy (anonymous):

determine integral of 2x + 1/ x^2 + x dx

OpenStudy (anonymous):

Let u=x^2+x Then du=2x+1 So your integral becomes: \[\int\limits \frac{du}{u}=\ln|u|+C=\ln|x^2+x|+C\]

OpenStudy (anonymous):

Should read: du=2x+1 dx***

OpenStudy (anonymous):

is that my final answer?

OpenStudy (anonymous):

Yessir :)

OpenStudy (anonymous):

how do you do integral 2y^3 e^2y4-2 dy

OpenStudy (anonymous):

Okay: \[\int\limits 2y^3 e^{2y^4-2}dy\] Let u=2y^4-2 du=8y^3dy so (1/4)du=2y^3dy So your integral becomes: \[\frac{1}{4}\int\limits e^u du\] What do you do from here?

OpenStudy (anonymous):

ok I got 1/4e ^(2y^4-1) + C

OpenStudy (anonymous):

Exactly :D

OpenStudy (anonymous):

Any other?

OpenStudy (anonymous):

yeah integral of xe^(2x squared) dx

OpenStudy (anonymous):

du=4x dx

OpenStudy (anonymous):

dx= 1/4 du

OpenStudy (anonymous):

We don't have to complicate, we just have to simplify the fuction that the integer give to us: \[\int\limits_{}^{}(((2x+1)\div(x^2))+x)dx\] \[2\int\limits_{}^{}(1/x) dx + \int\limits_{}^{}(1/x^2)dx + \int\limits_{}^{}xdx\] So, \[2\ln x-(1/x)+((x^2/2))+C\] I hope i help you